# Presentations of Schur and Specht modules in characteristic zero

### v1

@@ 5. Schur modules and Garnir relations
$\Lambda^{a}\otimes\Lambda^{k}\otimes\Lambda^{b-k}\xrightarrow{\gamma_{k}}\Lambda^{a}\otimes\Lambda^{b}\xrightarrow{\pi_{\mu}}L_{\mu},$
@@
(5.1) $B(a,b,k,i)=\{(t,i-t)\in\mathbb{N}\times\mathbb{N}:p\leq t\leq q\},$
@@
(5.2) $\pi_{\mu}\circ\delta_{i}\circ\gamma_{k}(x\otimes y\otimes z)=\sum_{j=p}^{q}\eta_{j}T(j)-\sum_{j=p}^{q}\epsilon_{j}\tbinom{a-k+j}{a-k}A(j),$
@@
(5.3) $\pi_{\mu}\circ\delta_{i}\circ\gamma_{k}(x\otimes y\otimes z)=\sum_{t=p}^{q}d_{k}(a,b,t)\eta_{t}T(t).$
@@
With the notation of the previous lemma, let $1\leq i\leq b$. (1) Suppose $a>k$. Let $p\neq 0$. By definition, $d_{k}(a,b,p)=1-(-1)^{p}\tbinom{a-k+p}{a-k}$. Since $a>k$ we have $\tbinom{a-k+p}{a-k}\neq 1$ and hence $d_{k}(a,b,p)\neq 0$.
@@
(5.4) $\operatorname{Im}(\overline{\gamma}_{k_{c}})\subseteq\operatorname{Im}(\Lambda^{\mu_{1}}\otimes\cdots\otimes\Lambda^{\mu_{c}+1}\otimes\Lambda^{\mu_{c+1}-1}\otimes\cdots\otimes\Lambda^{\mu_{\ell}}\xrightarrow{1\otimes\cdots\otimes\theta_{1}\otimes\cdots\otimes 1}\Lambda^{\mu}).$
@@
6. Presentations of Specht modules
@@
6.1. Specht modules and symmetrizations of Garnir relations
@@
Let $a,b,k\in\mathbb{N}$ such that $a\geq b\geq k\geq 1$. Then as $\mathfrak{S}_{a+b}$-modules $M^{(a,b)^{\prime}}/SGR^{(a,b)^{\prime}}(k)$ $\simeq S^{(a,b)^{\prime}}$ if and only if $\sum_{t=1}^{j}(-1)^{t-1}\tbinom{b-t}{b-k}\tbinom{j}{t}\tbinom{a-b+j+t}{t}\neq 0$ for all $j=1,\dots,k$.
@@
for all $c=1,\dots,\lambda_{1}-1$ and $j=1,\dots,k_{c}$.
@@
This follows by applying the Schur functor to Corollary 3.6 for the choice $\Psi_{\mu(c)}=\psi_{k_{c}}\in\operatorname{Hom}_{G}(\Lambda^{\mu_{c}}\otimes\Lambda^{\mu_{c+1}},\Lambda^{\mu_{c}}\otimes\Lambda^{\mu_{c+1}})$, $c=1,\dots,\lambda_{1}-1$. From (6.1) we have $\operatorname{Im}(f(\sum_{c=1}^{{\ell}-1}\overline{\psi}_{k_{c}}))=SGR^{\lambda}(k_{1},\dots,k_{\lambda_{1}-1})$. We know from the second equation of Lemma 4.3 that $a_{\mu(c),t}=(-1)^{t-1}$ $\tbinom{\mu_{c+1}-t}{\mu_{c+1}-k_{c}}$ if $t\neq 0$ and $a_{\mu(c),t}=0$ if $t=0$. ∎
@@
be the sum in Corollary 6.1, where $1\leq j\leq k$. (1) Let $k=1$. Then $\Sigma(1,1)=a-b+2$ and since $a\geq b$ we have $\Sigma(1,1)\neq 0$. From Theorem 6.2 we conclude that for any partition $\lambda$ we have
@@
Thus we recover Theorem 3.5 of [1] (which we have stated as Theorem 1.2 in the Introduction). (2) Let $k=b$ the maximal value of $k$. First we note that in this case there is no symmetrization, i.e. we have $SGR^{(a,b)^{\prime}}{(b)}=GR^{(a,b)^{\prime}}(b)$. Using the first identity of [7, Lemma 4.1(2)], we conclude that $\Sigma(b,j)=-(-1)^{j}\tbinom{a-b+j}{j}+1.$ Taking into account that $j\neq 0$, we see that the left hand side is equal to 00 if and only if $j>0$ is even and $a=b$. Since $j\leq b$, it follows from Corollary 6.1 that $M^{(a,b)^{\prime}}/GR^{(a,b)^{\prime}}(b)=S^{(a,b)^{\prime}}$ if and only if $b=1$ or $b>1$ and $a\neq b$. Hence from Theorem 6.2 we conclude that
@@
for any partition $\lambda$ satisfying $\lambda^{\prime}_{c}>\lambda^{\prime}_{c+1}$ for those $c$ for which $\lambda^{\prime}_{c+1}>1$. Thus we recover Theorem 1.1 of [3] (which we have stated as Theorem 1.1 in the Introduction). (3) Let $k=2$. Then $\Sigma(2,1)=(b-1)(a-b+2).$ Since $b\geq 2$ and $a\geq b$, we have $\Sigma(2,1)\neq 0$. A quick computation yields $\Sigma(2,2)=-1/2(a-b+3)(a-5b+8)$. Since $a\geq b$, we have $\Sigma(2,2)\neq 0$ if an only if $a\neq 5b-8$. From Corollary 6.1, $M^{(a,b)^{\prime}}/SGR^{(a,b)^{\prime}}{(2,2)}\simeq S^{(a,b)^{\prime}}$ if and only if $a\neq 5b-8.$ According to Theorem 6.2, if $\lambda$ is a partition with all column lengths greater than or equal to 2, then
@@
if $\lambda^{\prime}_{c}\neq 5\lambda^{\prime}_{c+1}-8$ for all $c=1,\dots,\lambda_{1}-1$.
@@
6.2. Specht modules and Garnir relations
@@
Consider the map $\gamma_{k}\in\operatorname{Hom}_{G}(\Lambda^{a}\otimes\Lambda^{k}\otimes\Lambda^{b-k},\Lambda^{a}\otimes\Lambda^{b})$ of Definition 5.1(2). It is straightforward to verify that its image $f(\gamma_{k})$ under the Schur functor $f$ has the property that the vector space $\operatorname{Im}(f(\gamma_{k}))\subseteq M^{(a,b)^{\prime}}$ is spanned by the Garnir relations $g_{1,k}^{T}$ of (1.1) as $T$ varies over $\mathcal{T}_{\lambda}$, where $\lambda=(a,b)^{\prime}$. Thus from Corollary 5.4 and Theorem 5.5 we obtain the following, which we regard as our third main result of the paper.
@@ Acknowledgment
We are indebted to Michelle Wachs for helpful comments concerning [3].

### v3

@@ 5 Schur modules and Garnir relations
$\Lambda^{a}\otimes\Lambda^{k}\otimes\Lambda^{b-k}\xrightarrow{\gamma_{k}}\Lambda^{a}\otimes\Lambda^{b}\xrightarrow{\pi_{\mu}}L_{\mu}$
@@
$B(a,b,k,i)=\{(t,i-t)\in\mathbb{N}\times\mathbb{N}:p\leq t\leq q\},$ (5.1)
@@
$\pi_{\mu}\circ\delta_{i}\circ\gamma_{k}(x\otimes y\otimes z)=\sum_{j=p}^{q}\eta_{j}T(j)-\sum_{j=p}^{q}\epsilon_{j}\tbinom{a-k+j}{a-k}A(j),$ (5.2)
@@
$\pi_{\mu}\circ\delta_{i}\circ\gamma_{k}(x\otimes y\otimes z)=\sum_{t=p}^{q}d_{k}(a,b,t)\eta_{t}T(t).$ (5.3)
@@
With the notation of the previous lemma, let $1\leq i\leq b$.
(1) Suppose $a>k$. Let $p\neq 0$. By definition, $d_{k}(a,b,p)=1-(-1)^{p}\tbinom{a-k+p}{a-k}$. Since $a>k$ we have $\tbinom{a-k+p}{a-k}\neq 1$ and hence $d_{k}(a,b,p)\neq 0$.
@@
$\operatorname{Im}(\overline{\gamma}_{k_{c}})\subseteq\operatorname{Im}(\Lambda^{\mu_{1}}\otimes\cdots\otimes\Lambda^{\mu_{c}+1}\otimes\Lambda^{\mu_{c+1}-1}\otimes\cdots\otimes\Lambda^{\mu_{\ell}}\xrightarrow{1\otimes\cdots\otimes\theta_{1}\otimes\cdots\otimes 1}\Lambda^{\mu}).$ (5.4)
@@
6 Presentations of Specht modules
@@
6.1 Specht modules and symmetrizations of Garnir relations
@@
Let $a,b,k\in\mathbb{N}$ such that $a\geq b\geq k\geq 1$. Then as $\mathfrak{S}_{a+b}$-modules $M^{(a,b)^{\prime}}/SGR^{(a,b)^{\prime}}(k)\simeq S^{(a,b)^{\prime}}$ if and only if $\sum_{t=1}^{j}(-1)^{t-1}\tbinom{b-t}{b-k}\tbinom{j}{t}\tbinom{a-b+j+t}{t}\neq 0$ for all $j=1,\dots,b$.
@@
for all $c=1,\dots,\lambda_{1}-1$ and $j=1,\dots,\mu_{c+1}$.
@@
This follows by applying the Schur functor to Corollary 3.12 for the choice $\Psi_{\mu(c)}=\psi_{k_{c}}\in\operatorname{Hom}_{G}(\Lambda^{\mu_{c}}\otimes\Lambda^{\mu_{c+1}},\Lambda^{\mu_{c}}\otimes\Lambda^{\mu_{c+1}})$, $c=1,\dots,\lambda_{1}-1$. From (6.1) we have $\operatorname{Im}(f(\sum_{c=1}^{{\ell}-1}\overline{\psi}_{k_{c}}))=SGR^{\lambda}(k_{1},\dots,k_{\lambda_{1}-1})$. We know from the second statement of Lemma 4.15 that $a_{\mu(c),t}=(-1)^{t-1}\tbinom{\mu_{c+1}-t}{\mu_{c+1}-k_{c}}$ if $t\neq 0$ and $a_{\mu(c),t}=0$ if $t=0$. ∎
@@
be the sum in Corollary 6.22, where $1\leq j\leq b$.
(1) Let $k=1$. Since $\tbinom{b-t}{b-k}=0$ if $t>1$, we have $\Sigma(1,j)=j(a-b+j+1)$ for every $j=1,\dots,b$. Since $a\geq b$, we have $\Sigma(1,j)\neq 0$ From Theorem 6.23 we conclude that for any partition $\lambda$ we have
@@
Thus we recover Theorem 3.5 of [1] (which we have stated as Theorem 1.2 in the Introduction).
(2) Let $k=b$ the maximal value of $k$. First we note that in this case there is no symmetrization, i.e. we have $SGR^{(a,b)^{\prime}}{(b)}=GR^{(a,b)^{\prime}}(b)$. Using the first identity of [8, Lemma 4.1(2)], we conclude that $\Sigma(b,j)=-(-1)^{j}\tbinom{a-b+j}{j}+1.$ Taking into account that $j\neq 0$, we see that the left hand side is equal to 00 if and only if $j>0$ is even and $a=b$. Since $j\leq b$, it follows from Corollary 6.22 that $M^{(a,b)^{\prime}}/GR^{(a,b)^{\prime}}(b)=S^{(a,b)^{\prime}}$ if and only if $b=1$ or $b>1$ and $a\neq b$. Hence from Theorem 6.23 we conclude that
@@
for any partition $\lambda$ satisfying $\lambda^{\prime}_{c}>\lambda^{\prime}_{c+1}$ for those $c$ for which $\lambda^{\prime}_{c+1}>1$. Thus we recover Theorem 1.1 of [3] (which we have stated as Theorem 1.1 in the Introduction).
(3) Let $k=2$. Then for every $j=1,\dots b$ we have
$\Sigma(2,j)=(b-1)j(a-b+j+1)-\tbinom{j}{2}\tbinom{a-b+j+2}{2}.$
A quick computation yields $\Sigma(2,j)=0$ if and only if $04(b-1)-(j-1)(a-b+j+2)=$ 0. From Corollary 6.22, $M^{(a,b)^{\prime}}/SGR^{(a,b)^{\prime}}{(2,2)}\simeq S^{(a,b)^{\prime}}$ if and only if $4(b-1)-(j-1)(a-b+j+2)\neq 0$ for all $j=1,\dots,b$. According to Theorem 6.23, if $\lambda$ is a partition with all column lengths greater than or equal to 2, then
@@
if $4(\lambda^{\prime}_{c+1}-1)-(j-1)(\lambda^{\prime}_{c}-\lambda^{\prime}_{c+1}+j+2)\neq 0$ for all $c=1,\dots,\lambda_{1}-1$ and $j=1,\dots,\lambda^{\prime}_{c+1}$.
@@
6.2 Specht modules and Garnir relations
@@
Consider the map $\gamma_{k}\in\operatorname{Hom}_{G}(\Lambda^{a}\otimes\Lambda^{k}\otimes\Lambda^{b-k},\Lambda^{a}\otimes\Lambda^{b})$ of Definition 5.17(2). It is straightforward to verify that its image $f(\gamma_{k})$ under the Schur functor $f$ has the property that the vector space $\operatorname{Im}(f(\gamma_{k}))\subseteq M^{(a,b)^{\prime}}$ is spanned by the Garnir relations $g_{1,k}^{T}$ of (1.1) as $T$ varies over $\mathcal{T}_{\lambda}$, where $\lambda=(a,b)^{\prime}$. Thus from and Corollary 5.20 and Theorem 5.21 we obtain the following, which we regard as our third main result of the paper.
@@ Acknowledgment
We are indebted to Michelle Wachs for helpful comments concerning [3]. Also, we would like to thank the referee for helpful suggestions.